### v4

# CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents

Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 96.4%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
[…]
tcb@breakable
CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents
Jing Gongthanks: Equal Contribution, Yanghui Wu11footnotemark: 1, Linxi Liang11footnotemark: 1, Yanlin Wangthanks: Corresponding Author, Jiachi Chen, Mingwei Liu, Zibin Zheng Sun Yat-sen University {gonj39,wuyh223,lianglx26}@mail2.sysu.edu.cn, yanlin-wang@outlook.com, {chenjch86,liumw26,zhzibin}@mail.sysu.edu.cn,
Introduction
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al., 2013; Sim et al., 2011; Stolee et al., 2014). Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al., 2020). Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR)  (Husain et al., 2019; Huang et al., 2021; Li et al., 2024a; Lu et al., 2021; Yao et al., 2018; Feng et al., 2020). According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average. For each query, they typically reference 2.64 code examples to find the necessary information. These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
Existing code search datasets like CoSQA  (Huang et al., 2021) and CodeSearchNet  (Husain et al., 2019) rely on expert human annotators to evaluate query-code pairs. While this approach captures human judgment, it exists several limitations. First, human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations. Second, manual annotation requires significant time and resources, limiting dataset scalability. Recent works by  Li et al. (2024a) leverage LLMs for annotation to address scalability issues. However, LLMs face challenges in reliably determining whether code functionally fully satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. Furthermore, most code search benchmarks employ metrics like MRR or Pass@k (Lu et al., 2021), which primarily evaluate single-result retrieval performance. This approach overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations. While alternative metrics like MAP and Recall@k Husain et al. (2019) exist for evaluating multiple code search, they underemphasize the importance of top-ranked results. This is problematic since, according to our survey, users heavily rely on first-page search results.
In this work, we begin with a survey of 102 experienced Python programmers to understand multi-choice code search practices. The survey results reveal that developers frequently require and utilize multiple code examples for a query. Motivated by these insights, we present a new benchmark CoSQA+, with 373,339 pairs of web queries filtered from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search. To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a first screener, test program generator, test executor, bug fixer, and final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks  (Wang et al., 2013). We use it to measure 9 code search methods, 8 of which are embedding-based.
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 200 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 96.4%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we develop the first Multiagent annotator with over 96.4% accuracy for code search dataset.
  • To the best of our knowledge, we construct CoSQA+ which is the first benchmark designed for evaluating multi-choice code search.
  • We enhance the fully automated construction process through candidate pair construction by multiple models and annotation by agents.
  • We publicly release CoSQA+ to facilitate further research in semantic code search.
Related Work
Code Search Datasets
Besides our work, there are other datasets and benchmarks in the realm of code search.  Husain et al. (2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub. CoSQA  (Huang et al., 2021) and CS1QA  (Lee et al., 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size. StaQC  (Yao et al., 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality.  Li et al. (2024b) introduces ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control. This benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague. Recently,  Li et al. (2024a) present Query4Code, which also leverages LLMs for dataset annotation, but their approach uses a single model for similarity computation and adopts a less strict matching criterion.
Test-Driven Agents
Traditional test generation approaches relied heavily on human expertise, but recent work has demonstrated the potential of automated test agents. LLM-based agents have shown promising results in various testing scenarios. For instance, in the domain of deep learning libraries,  Deng et al. (2024) develops agents capable of generating diverse and meaningful test inputs. Several studies have explored LLM-based agents for specialized tasks. For test oracle generation,  Nashid et al. (2023) utilizes few-shot prompting to create context-aware assertions, achieving a high accuracy. In program repair, studies by  Xia and Zhang (2023) and  Jiang et al. (2023) have demonstrated the effectiveness of integrating LLMs with iterative validation workflows. However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance.
Figure 1: The construction process of CoSQA+. The final CoSQA+ dataset includes queries from original CoSQA, and codes filtered from CodeSearchNet.
CoSQA+
Figure 1 illustrates the benchmark construction process for CoSQA+. The process starts with query and code collection, in which we leverage the queries from CoSQA and collect python code snippets filtered from CodeSearchNet. In candidate pairs construction, collected code snippets and queries are embedded as vectors to calculate similarity. For each query, the top 20 query-code pairs with the highest similarity are selected as candidates for potential matches. In model annotation, each query-code pair is annotated by test-driven agents with “1” or “0”, indicating whether the code snippet exactly matches the query. These labeled pairs are then added to CoSQA+.
Query and Code Collection
This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus  (Husain et al., 2019) serves as the code sources.
Query Source
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes  (Huang et al., 2021). We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets  (Sun et al., 2022). (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent. This careful curation helps maintain high data quality and reduces noise in the training process  (Li et al., 2020). By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
Code Source
To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages  (Husain et al., 2019). According to our survey, code searchers prefer function codes with clear input and output. These codes are high quality with higher readability, testability and maintainability  (Martin, 2009). To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach. This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement. Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
Candidate Query-Code Pairs Construction
Labeling all query-code pairs is impractical. To improve efficiency, we only retain high-confidence pairs before annotation. This section details the construction process of high-confidence query-code pairs.
Construction Steps
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. For a detailed breakdown of query distributions by the number of exactly matched codes, please see Appendix D. Distribution of Queries Based on Exactly Matched Codes.
Multi-model Embedding
High similarity between query and code is fundamental for a high-quality dataset  (Sun et al., 2022). During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. And this will potentially lead to unfair test result of benchmark. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. Specifically, we use three models: CodeBERT  (Feng et al., 2020), UniXcoder  (Guo et al., 2022) and CodeT5+ 110M embedding  (Wang et al., 2023).
The relatedness between the query and code is quantified by the cosine similarity, calculated as:
${cosine\_similarity}=\frac{\mathbf{v}_{\text{query}}\cdot\mathbf{v}_{\text{code}}}{\|\mathbf{v}_{\text{query}}\|\|\mathbf{v}_{\text{code}}\|}$
where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
Top-20 Code Candidates Selection
Before constructing the full set of candidate pairs, we randomly selected 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query. The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Figure 2). As shown in our analysis, when the number of candidates reaches 20, the marginal benefit of further expansion diminishes. This decline can be attributed to limitations of codebase, nature of queries and ability of embedding models. Furthermore, adding just one more candidate code per query results in an additional 20,604 pairs, which would incur substantial costs in terms of both time and financial resources for annotation. Thus, limiting candidates to the top-20 strikes a balance between maximizing the retrieval of relevant codes and controlling annotation costs.
Figure 2: Probability of identifying matched codes as the number of candidate codes increases
Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design a test-driven annotation agents based on DeepSeek-V3. An preliminary screener will determine whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation. All prompts and a detailed example are provided in AppendixA.
Figure 3: Example illustration of (1) clear match, (2) clear non-match, and (3) unclear cases.
Preliminary Screener
The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Figure 3. Guided by the prompt instructions, the model assesses whether the code fully addresses the query’s functional requirements. For unambiguous cases, it provides a justification and immediate classification. For ambiguous cases (representing 33% of candidates), it flags the pair for deeper verification by downstream agents. This triage mechanism optimizes the annotation workflow by resolving straightforward cases upfront while reserving computational resources for ambiguous scenarios.
Test Program Generator
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, the test program is generated based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation.
Test Executor and Bug Fixer
The test executor operates within an isolated Docker environment to ensure safe execution while preventing host system interference. When errors such as ModuleNotFoundError occur, the pipeline initiates a two-stage repair process. First, it attempts direct installation of missing dependencies using standard package managers. If unresolved, the system invokes an LLM to generate context-aware terminal commands for dependency resolution, which are executed iteratively until the error is mitigated or deemed irreparable. After dependency fixes, the test suite is re-executed to validate functionality. Execution traces, assertion outcomes, and error logs are aggregated into a structured report, which is forwarded to the final arbiter for assessment. This approach enhances the reliable verification of ambiguous cases while maintaining environmental integrity, achieving an executable program ratio of 83.67%.
Final Arbiter
The framework ultimately concludes with a final arbiter stage, where the agent employs a specialized prompt to assess if the generated code fully satisfies the query. This assessment is based on both the test program produced and the outcomes of its execution. This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior. By grounding judgments in dynamic test execution rather than relying exclusively on static structural analysis and semantic analysis. An example is provided in Appendix BB. An example of Final Arbiter.
Benchmark Statistics
In total, 412,080 labeled query-code pairs are retained. Table 1 shows the statistics of our CoSQA+ benchmark.
Number of pairs Number of code Code average length Average test cases Executable rate
412,080 132,952 378.052 4.31 83.67%
Table 1: Statistics of CoSQA+.
Cost Analysis
For data annotation, a total of 412,080 data samples are labeled by Deepseek-V3 at a cost of approximately \$257, resulting in an average cost of \$0.00062 per sample. The labeling time varies significantly based on complexity: clear cases require approximately 10 seconds per sample, while non-clear cases may take up to 50 seconds, depending on API rate limits, dependency installations, and network latency. However, the annotation process benefits from full concurrency across multiple queries, enabling rapid completion when computational resources are sufficient.
Evaluation
In this section, we conduct experiments with corresponding analysis to answer the following research questions (RQs):
RQ1: Is the quality of CoSQA+ higher than that of CoSQA?
RQ2: Can our test-driven agents surpass Python experts and LLMs in the annotation task?
RQ3: How effective are the existing approaches in multi-choice code search on our benchmark?
RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
Metrics
Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG  (Wang et al., 2013) and MRR  (Feng et al., 2020). NDCG measures the quality of ranking by considering both the relevance and position of retrieved results. For a given query $q_{i}$, we calculate NDCG@k as:
$NDCG@k=\frac{DCG@k}{IDCG@k}$
where DCG@k is the Discounted Cumulative Gain up to position k, and IDCG@k is the ideal DCG@k (the maximum possible DCG@k). The final NDCG score is averaged across all queries. MRR evaluates the average reciprocal of the rank at which the first relevant code appears:
$MRR=\frac{1}{|N|}\sum_{i=1}^{|N|}\frac{1}{rank_{i}}$
where |N| is the number of queries and rank_i is the rank of the first relevant code for the i-th query.
Experimental Setup
In the first experiment, we randomly select 1000 query-code pairs labeled "1" from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, so we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
For the second experiment, we fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with NDCG@10 and MRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
Model Metric Fine-tune Dataset
None CoSQA CoSQA+
CodeBERT MRR 0.952 0.939 0.966
NDCG@10 0.958 0.937 0.962
UniXcoder MRR 0.891 0.951 0.954
NDCG@10 0.890 0.946 0.952
CodeT5+ MRR 0.794 0.938 0.940
embedding NDCG@10 0.806 0.943 0.951
Table 2: Performance of code search models fine-tuned with different data. The test set is CSN Python.
Results and Analysis
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 73.8% of the better codes are from CoSQA+, while 26.2% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table 2, the three models fine-tuned with CoSQA+ outperforms those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
RQ2: Performance of Agent-based Annotator
We design this experiment to assess whether our agent-based annotation approach can achieve comparable or superior quality to traditional human annotation and single LLM annotation.
Experimental Setup
To establish a reliable ground truth, we first have three Python experts design test programs for 300 randomly selected query-code pairs and annotate them based on test execution results. The final ground truth labels are determined by majority voting among these three experts. We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators.
Results and Analysis
Method Accuracy
Agent-based Annotator 0.964
Human Wo. Test Program 0.890
Deepseek-V3 0.820
O1-mini 0.816
Claude-3.5-Sonnet 0.830
Table 3: Accuracy comparison of different annotation methods against ground truth.
As shown in Table 3, our agent-based annotator achieves 96.4% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (82.0%), O1-mini (81.6%) and Claude-3.5-Sonnet (83.0%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
Importantly, about 83% of the test programs generated by our agent are successfully executable, indicating the reliability of our test program generation approach. The high executability rate indicates that most query-code pairs can be verified through actual program testing rather than relying solely on semantic analysis.
The superior performance of our agent-based annotators can be attributed to two key factors: 1) The test program provides concrete evidence of whether the code satisfies the query requirements, reducing ambiguity in the annotation process; 2) The systematic verification through test execution helps eliminate human bias and inconsistency that often occur in manual annotation.
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.85, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 96.4% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
RQ3: Existing Methods Performance on CoSQA+
This experiment aims to evaluate the effectiveness of different models in accurately retrieving code snippets that match natural language queries on the CoSQA+ benchmark.
Task
The natural language code search task is a text retrieval problem. Given a natural language query $q_{i}$ and a collection of code snippets $C=\{c_{1},\ldots,c_{n}\}$, the goal is to identify all code snippets
$\{c^{*}_{j}\mid c^{*}_{j}\in C\text{ and }c^{*}_{j}\text{ exactly matches }q_{i}\}$
that exactly match the query.
Experimental Setup
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric. NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly.
The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach (Zhang et al., 2010) with state-of-the-art semantic code search models, namely: CodeBERT (Feng et al., 2020), UniXcoder (Guo et al., 2022), and CodeT5+ embedding (Wang et al., 2023). In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3  (Sturua et al., 2024), All-MiniLM-L12-v2 (Reimers and Gurevych, 2021), All-mpnet-base-v2  (Song et al., 2021), and Multilingual-e5-large  (mul, 2023).
Results and Analysis
Method NDCG@10 MRR
BoW 0.046 0.034
CodeBERT 0.232 0.175
UniXcoder 0.187 0.140
CodeT5+ embedding 0.179 0.136
Jina-embedding-v3 0.225 0.171
All-MiniLM-L12-v2 0.214 0.163
All-mpnet-base-v2 0.227 0.172
Multilingual-e5-large 0.218 0.166
Table 4: Performance of Various Methods on CoSQA+ Using NDCG@10 and MRR Metrics.
While traditional information retrieval methods, such as BoW, are computationally efficient, Table 4 demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist. Among the deep learning-based models, CodeBERT achieves the highest performance with an NDCG@10 of 0.232 and an MRR of 0.175. This strong performance can be attributed to its encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances its code understanding capabilities. Other embedding-based models also show substantial improvements over traditional methods, highlighting the benefits of modern deep learning architectures for semantic code search.
RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search. Notably, CodeBERT leads the performance ranking by achieving the highest scores in both NDCG@10 and MRR, while the alternative embedding-based models also offer competitive results.
Conclusion
In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA. The test-driven annotation approach achieves a high accuracy of 96.4%, outperforming traditional human annotation and single LLM methods. Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks. We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task
Ethical consideration
In this work, we prioritize ethical considerations by the responsible use of data and methodologies. The dataset sources, CoSQA and CodeSearchNet, are publicly available under permissible licenses, mitigating privacy concerns. Our test-driven agents are designed to minimize human bias by automating code evaluation, thus enhancing fairness and consistency. Additionally, we openly release CoSQA+ to promote transparency and reproducibility in code search research. While our approach aims to improve code search efficiency, we acknowledge the potential for misuse in generating or retrieving malicious code. We encourage the community to adopt ethical guidelines when using our benchmark and tools, ensuring that their application aligns with positive social impact.
Limitations
We are committed to using LLM-based test-driven annotation to construct high-quality function-level multi-choice code search datasets, but there are still several limitations. Firstly, due to resource constraints, we only apply the construction method to Python code. The method is generalizable, and we plan to extend it to other programming languages. Secondly, we conduct extensive experiments, research, and manual validation to demonstrate the effectiveness of the construction method and the quality of the dataset, but due to limitations in time and resources, CoSQA+ may contain some biases or ignore some scenarios. Finally, although various mechanisms are designed to improve the success rate and accuracy of automated annotation and achieve promising results, factors such as ambiguity in query statements, complexity of code functionality, and uncertainty in program execution environments prevent ideal outcomes for few query-code pair annotations. We plan to continue exploring and improving the framework and methods for automated annotation in the future.
Appendix A A. Prompts for Test-Driven Agents
Preliminary Screener
1def get_preliminary_screening_prompt( query, function_code):
2 prompt = f"""
3 Instructions for Evaluation:
4 You are provided with a query that specifies requirements for a function and a block of function code.
5 Your goal is to determine if the code fully matches the query’s requirements.
6 - If the code clearly does not satisfy the query, output preliminary_screening: 0 and provide a corresponding explanation.
7 - If the code clearly satisfies the query, output preliminary_screening: 1 and provide a corresponding explanation.
8 - If the determination is ambiguous or if further testing using a test_program_generation agent is needed, output preliminary_screening: 0.5 and provide a corresponding explanation.
9
10 Ensure your answer is formatted exactly as follows:
11 preliminary_screening: [0/0.5/1],
12 explanation: <Your explanation in English>
13
14
[…]
17 """
18 return prompt
Test Program Generator
1def get_test_program_prompt(query, function_code):
2 prompt = f"""
3 Task Description:
4 Please generate a test program based on the provided query and function code. The test program should:
5 1. Contain assert statements in the main function to verify that the function code correctly implements the behavior described in the query.
6 2. **Do not** test other classes or functions generated, only focus on testing the behavior specified in the query.
7 3. The generated test program should be complete, executable, and must test the functionality described by the query using assert statements.
8
9 For example, given the following query and function code:
10
11 query: "python remove all empty items in list"
12 code:
13 def remove_from_list(list, remove_items):
14 list = return_list(list)
15 remove_items = return_list(remove_items)
16 for remove in remove_items:
17 if remove in list:
18 list.remove(remove)
19 return list
20
21 The generated program should:
22 Test the behavior described in the query (removing empty items from a list).
23 Include only the assert statements in the main function.
24 Be a fully executable test program that can verify the behavior of function code.
25
26 Example Test Program Output:
27
28 # Additional function definition
29 def return_list(input_list):
30 return [item for item in input_list if item]
31
32 # Original function code
33 def remove_from_list(list, remove_items):
34 list = return_list(list)
35 remove_items = return_list(remove_items)
36 for remove in remove_items:
37 if remove in list:
38 list.remove(remove)
39 return list
40
41 # Test code
42 def main():
43 # Test case 1: Remove empty items from the list
44 list1 = [1, 2, "", 4, None, "test", ""]
45 remove_items1 = [""]
46 result1 = remove_from_list(list1, remove_items1)
47 # Assert that empty strings have been removed from the list
48 assert result1 == [1, 2, 4, None, "test"]
49
50 # Test case 2: Remove None from the list
51 list2 = [1, 2, "", 4, None, "test", ""]
52 remove_items2 = [None]
53 result2 = remove_from_list(list2, remove_items2)
54 # Assert that None has been removed from the list
55 assert result2 == [1, 2, "", 4, "test", ""]
56
57 # Test case 3: Ensure no changes when removing an item not in the list
58 list3 = [1, 2, 3]
59 remove_items3 = [4]
60 result3 = remove_from_list(list3, remove_items3)
61 # Assert that the list remains unchanged
62 assert result3 == [1, 2, 3]
63
64 # Test case 4: Empty list case
65 list4 = []
66 remove_items4 = [""]
67 result4 = remove_from_list(list4, remove_items4)
68 # Assert that the result is still an empty list
69 assert result4 == []
70
71 print("All tests passed.")
72
73 if __name__ == "__main__":
74 main()
75
76 Now please directly and only generate the complete test program based on the instructions above. Do NOT add any explanation.
77 query: {query}
78 code: {function_code}
79 """
80
81 return prompt
Bug Fixer
1def get_pip_install_command_prompt( test_output: str, test_program: str) -> str:
2 prompt=f"""You are a Python development assistant. The user encountered a ModuleNotFoundError when running their test code. Here are the details:
3 Test Output:
4 {test_output}
5
6 Test Program:
7 {test_program}
8
9 Please determine the correct pip install command(s) required to install the missing module(s). Consider that the PyPI package name might differ from the module name. Provide only the necessary pip install command(s), each on a separate line. Do not include any other text or explanation.
10
11 Examples of valid responses:
12 1. If missing ’cv2’:
13 pip install opencv-python
14
15 2. If missing ’yaml’:
16 pip install pyyaml
17
18 3. If multiple packages needed:
19 pip install numpy
20 pip install pandas
21
22 Now, provide the correct pip install command(s) for the current error:
23 """
24 return prompt
Final Arbiter
1 def get_final_judge_prompt( function_code: str, query: str, test_program: str, test_result: str,
2 test_output: str) -> str:
3 prompt = f"""
4 You are an code expert in judging whether a function code can fully satisfy the requirement of the query. You can refer to a generated test program and its test result and test output.
5
6 You are provided with the following inputs:
7
8 1. Query:
9 {query}
10
11 2. Function Code:
12 {function_code}
13
14 3. Test Program:
15 {test_program}
16
17 4. Test Result:
18 {test_result}
19
20 5. Test Output:
21 {test_output}
22
23 Your Task:
24 Evaluate whether the provided function code completely satisfies the query’s requirements based on a detailed analysis of the function code and the results and the output from running the test program.
25
26 Judgment Guidelines:
27 - If the function code clearly fails to meet the query requirements (supported by incorrect logic or failing test results), then output:
28 final_verdict: 0,
29 final_explanation: "A clear explanation describing why the function code does not meet the query requirements, referencing any issues in the code and test failures."
30
31 - If the function code fully meets all the query requirements (supported by correct logic and successful test results), then output:
32 final_verdict: 1,
33 final_explanation: "A clear explanation describing how the function code meets all the query requirements, referencing the successful test outcomes and correct implementation."
34
35 Important:
36 - Your final output must strictly adhere to the following format:
37 final_verdict: [0 or 1],
38 final_explanation: <Your explanation in English>
39 - Only output 0 or 1 as the final verdict.
40
41 Please provide your final judgment as instructed.
42 """
43 return prompt
B. An example of Final Arbiter
This figure illustrates the process of handling an unclear case. The test program generator generates a program, which is then executed by the test executor. The test results are returned to the final arbiter, which assigns a label to the case.
Figure 4: Process of Handling an Unclear Case through Test Program Generation, Execution, and Final Arbitration
C. Code Search Behavior Survey
We conduct this survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.
Survey Demographics and Methodology
A total of 102 participants complete the survey, predominantly undergraduate students (89.22%) with some master’s students (8.82%). The majority of participants (92.16%) have formal computer science education, and most have 1-5 years of programming experience (66.67%), while 30.39% have less than one year of experience. All participants are Chinese university students who use Python for development, with 52.94% considering Python as their primary programming language. The survey is conducted anonymously through an online platform, with all questions presented in Chinese. No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
Survey Results and Analysis
Our survey reveal several significant patterns in code search behavior. Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution. Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%). When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
Survey Questions
Table 5 shows the key questions from our survey (translated from Chinese).
Questions Response Rate (%)
A B C D
How often do you perform code searches during programming-related tasks per day?
(A: >10 times, B: 6-10, C: 3-5, D: 1-2) 50.00 19.61 14.71 15.68
In your daily development, how many code examples do you typically need to reference for a query (such as API usage, algorithm implementation)
(A: Only 1, B: 2-3, C: 4-5, D: >5) 16.67 66.67 12.75 3.92
In which scenarios do you most frequently need to reference multiple code examples when performing code searches? [Multiple Choice]
(A: API usage, B: Algorithm implementation, C: Performance optimization solutions, D: Other) 68.63 60.78 51.96 47.06
When performing code searches, which better matches your habits?
(A: Only need to find one code snippet that meets the requirements, B: Need to find more than one snippet that meets the requirements) 49.02 50.98 - -
When search returns multiple code examples, which aspects do you focus on most? [Multiple Choice]
(A: Differences between code examples, e.g. comparison between different implementation methods, B: Relevance of each example, e.g. degree of matching with the query, C: Whether the display order of examples meets expectations, e.g. most relevant first, D: Whether examples cover common use cases or edge conditions) 56.86 65.69 49.02 38.24
When evaluating code search results, which do you prioritize more?
(A: Position of the first result that meets requirements, B: Average position of the first few results that meet requirements, C: Diversity of results, D: Other) 44.12 37.25 18.63 0.00
When searching for code examples, which type of function structure do you prefer?
(A: Functions with both input and output, B: Functions with input but no output, C: Functions with output but no input , D: Functions with neither input nor output) 71.57 16.67 8.82 2.94
When searching for code examples, which type of function structure do you end up using?
(A: Functions with both input and output, B: Functions with input but no output, C: Functions with output but no input , D: Functions with neither input nor output) 62.75 22.55 11.76 2.94
Table 5: Distribution of survey responses (n=102). For each question, the options are listed in parentheses.
D. Distribution of Queries Based on Exactly Matched Codes
The table 6 illustrates the distribution of queries according to the number of exactly matched codes, providing insight into the frequency of queries with varying code match counts.
Table 6: Distribution of Queries Based on Exactly Matched Codes
Number of Exactly Matched Codes (n) Query Count
1 2907
2 2445
3 2121
4 1748
5 1429
6 1266
7 1108
8 873
9 724
10 628
11 481
12 398
13 345
14 220
15 155
16 126
17 78
18 44
19 19
20 3