### v5

# CoSQA+: Enhancing Code Search Evaluation with a Multi-Choice Benchmark and Test-Driven Agents

Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 92.0%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
Index Terms:
software engineering, information search and retrieval, human-computer interaction.
11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
I Introduction
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development [1, 2, 3, 4, 5]. Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems [6]. Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12]. According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average. For each query, they typically reference 2.64 code examples to find the necessary information. These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
However, existing approaches face significant challenges in scalability, annotation quality, and evaluation metrics:
  1. C1
    Scalability Issues in Human-Based Benchmarking. Conventional benchmarks like CoSQA [8] and CodeSearchNet (CSN) [7] rely on expert human annotators to evaluate query-code pairs. While this approach captures human judgment, manual annotation requires significant time and resources, limiting dataset scalability. The labor-intensive nature of this process makes it impractical for creating large-scale benchmarks needed for robust code search evaluation.
  2. C2
    Accuracy Limitations in Annotation Methods. Both human and LLM-based annotation approaches suffer from accuracy issues. Human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations. While Li et al. leverage LLMs to address scalability issues [9], these models struggle to reliably determine whether code functionally satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. This fundamental limitation affects the quality of both human and LLM-based annotations [13].
  3. C3
    Inadequate Evaluation Metrics for Code Search. Most code search benchmarks employ metrics like MRR or Pass@k [10] which primarily evaluate single-result retrieval performance. This overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations. While alternative metrics like MAP and Recall@k [7] exist for evaluating multiple code matches, they underemphasize the importance of top-ranked results. This is particularly problematic as our survey reveals users heavily rely on first-page search results, highlighting the need for better evaluation metrics.
To address these limitations, we present a new benchmark CoSQA+, with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search. To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a preliminary screener, a test program generator, a test executor, a bug fixer, and a final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks [14]. We use it to measure 8 code search methods, 7 of which are embedding-based.
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 400 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and a single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 92.0%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we develop the first Multiagent annotator with 92.0% accuracy for code search dataset.
  • To the best of our knowledge, we construct CoSQA+ which is the first benchmark designed for evaluating multi-choice code search.
  • We enhance the fully automated construction process through candidate pair construction by multiple models and annotation by agents.
  • We publicly release CoSQA+ to facilitate further research in semantic code search.
II Motivation Survey
Do developers really need multi-choice code search? To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.Fig. 1 provides a multi-choice code search example
Figure 1: An example of multi-choice code search. The query is vague in its definition of "illegal character," allowing eight candidate code snippets to match under different interpretations.
Methodology
The survey targets 102 participants, predominantly undergraduate students (89.22%) with a smaller proportion of master’s students (8.82%). Most of the participants (92.16%) have formal computer science education, and their experience in Python programming varies. The majority (66.67%) have between 1 and 5 years of programming experience, while 30.39% have less than one year of experience. The participants are selected from programmers using Python as their primary programming language (52.94%). The survey is conducted anonymously through an online platform. No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
Results and Analysis
Our survey reveals several significant patterns in code search behavior. Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution. This suggests that the current code search task, which typically prioritizes a single result, may need to evolve to provide multiple relevant code snippets. Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%). These preferences emphasize that developers do not just want the closest match to their query, but also value diverse approaches that help them understand the problem and solution better. When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
III Related Work
III-A Code Search Datasets
Besides our work, there are other text-to-code datasets and benchmarks in the realm of code search. They can be categorized into three types based on the origin of the queries: doc-code datasets, question-code datasets, and query-code datasets.
CodeSearchNet [7], CoDesc [15], PyMT5 [16], and XLCoST [17] are doc-code datasets, which contain massive code snippets from public repositories with corresponding comments. They lay foundational training grounds for code search models to grasp basic code-comment relationships. CodeSearchNet [7] dataset comprises a large-scale collection of query-code pairs extracted from GitHub repositories. CoDesc [15] is a comprehensive dataset featuring 4.2 million Java methods paired with natural language descriptions, meticulously filtered to minimize noise. PyMT5, on the other hand, offers a vast collection of 26 million Python methods and 7.7 million method-docstring pairs, enabling bidirectional translation between code and docstrings. XLCoST [17] stands out as the largest multilingual parallel dataset, encompassing 8 languages. It supports 10 cross-lingual code tasks, such as translation and search, and provides essential baselines for model development in diverse programming contexts.
StaQC [11], CoNaLa [18], SO-DS [19], CS1QA [20], and ProCQA [21] are question-code datasets, which contain code answers with corresponding questions from Stack Overflow. They challenge models to excel in practical problem-solving scenarios by understanding complex developer inquiries. StaQC [11] is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality. CoNaLa [18] is a curated dataset derived from Stack Overflow, comprising 2,379 training and 500 test examples of natural language-code pairs, alongside a large automatically-mined dataset of 600k examples, designed for tasks like code synthesis, retrieval, and summarization. SO-DS [19], a high-quality snippet collection mined from highly upvoted Stack Overflow posts, focuses on annotated code search, offering 12,137 snippets and 7,674 unique descriptions, with 2,225 annotated queries for validation and testing, specifically tailored for improving code retrieval systems by leveraging natural language descriptions. CS1QA [20], a dataset for code-based question answering in programming education, comprises 9,237 annotated question-answer pairs from Python chat logs, alongside 17,698 unannotated examples, designed to benchmark models on code comprehension and question answering in an educational context. ProCQA [21] introduces a large-scale community-based programming question answering dataset, which is mined from StackOverflow with strict filtering strategies for quality and fairness control. This benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague.
CoSQA [8], CSN99 of CodeSearchNet [7], WebQueryTest of CodeXGLUE [10], and Query4Code [9] are query-code datasets, which contain real human queries from web and code snippets matched with them. They are crucial for training and testing models’ capabilities in handling authentic, diverse search queries. CoSQA [8] introduces a dataset of 20,604 human-annotated natural language query-code pairs, enhancing semantic matching for code search, and proposes CoCLR, a contrastive learning method that boosts code question answering accuracy by 10.5% when combined with CodeBERT. CSN99, part of the CodeSearchNet Challenge [7], offers 99 natural language queries with approximately 4k expert-annotated relevance judgments, facilitating robust evaluation of code search models. The queries of CoSQA and CSN99 are collected from Microsoft Bing search engine. WebQueryTest [10] addresses the gap between real user search queries and existing code search datasets by providing a testing set of about 1k real web queries paired with Python code functions, classified for code-search intent and query-code relevance, with annotations from 13 proficient developers. Query4Code leverages Large Language Models (LLMs) to annotate datasets, primarily relying on the semantic information embedded within the code [9]. However, it lacks test cases or validation mechanisms to ensure the accuracy and reliability of the annotations.
Based on these datasets, COIR [22] emerges as a robust and comprehensive benchmark specifically designed for code retrieval. COIR comprises ten curated datasets spanning eight retrieval tasks across seven diverse domains, offering a versatile evaluation platform for state-of-the-art retrieval models. Its user-friendly Python framework, compatible with popular benchmarks like MTEB [23] and BEIR [24], facilitates seamless integration into research workflows.
Despite the advancements of these code search datasets, several limitations persist. Doc-code datasets oversimplify the search task by focusing on direct comment-to-code mappings, lacking the complexity of real-world queries. Question-code datasets are not designed for code search tasks, emphasizing question-answer context over direct code retrieval. Query-code datasets, though highly relevant, are constrained by their difficulty in constructing and annotating at scale.
III-B Code Search Methods
The existing methods for code search can mainly be divided into two mainstreams: information retrieval (IR) based methods and deep learning (DL) based models. IR-based methods extract keywords from queries and search for these keywords in the code base to retrieve the most relevant code snippets [25, 26, 27]. However, standard information retrieval methods have drawbacks in the code search domain, as they treat codes as text while programming languages and natural languages differ greatly. Fortunately, DL-based methods receive great attention in NLP and achieve significant breakthroughs in code search research. The CodeSearchNet [7] dataset paves the way for CodeBERT [12] and GraphCodeBERT [28], both of which are bimodal pre-trained BERT [29] based models that leverage Masked Language Modeling (MLM) and Replaced Token Detection (RTD) [30] objective. UniXcoder [31] unifies three pre-training designs into one architecture and utilizes AST structure and code comment to enhance the cross-modal alignment. Recent T5 [32] based models like CodeT5 [33] and CodeT5+ [34] show outstanding performance in code search, utilizing the text-to-text transformer architecture and pre-training on large code datasets. While the two models achieve high performances in the CSN benchmark, GraphCodeBERT [28] demonstrates superior results by harnessing data flow encoded in the Abstract Syntax Tree (AST) of code to enrich the structural information during pre-training. Recent advancements in text embedding models have significantly enhanced the performance of code search methods, particularly in multilingual and long-context retrieval tasks. Among these, jina-embeddings-v3 [35] stands out as a state-of-the-art model with 570 million parameters, achieving superior performance on the MTEB benchmark. Its innovative use of task-specific Low-Rank Adaptation (LoRA) [36] adapters enables high-quality embeddings for various tasks, including query-document retrieval and text matching, while supporting context lengths of up to 8192 tokens. This model outperforms both proprietary embeddings like OpenAI and Cohere. multilingual-e5-large [37, 38], which is trained on 1 billion multilingual text pairs and fine-tuned on labeled datasets, is efficient, especially in reducing embedding dimensions without compromising quality, thanks to Matryoshka Representation Learning [39]. But it does not match the flexibility and performance of jina-embeddings-v3. On the other hand, all-MiniLM-L12-v2 [40] and all-mpnet-base-v2 [41] represent efficient, compressed models derived from larger pre-trained Transformers. The all-MiniLM-L12-v2 model, based on MiniLM [42], employs deep self-attention distillation to retain high accuracy with reduced parameters, making it suitable for latency-sensitive applications. Similarly, all-mpnet-base-v2, built on MPNet [43], combines the strengths of BERT and XLNet [44] by leveraging masked and permuted language modeling, achieving competitive results on downstream tasks [45]. While these models are effective for general-purpose embeddings, they lack the multilingual and long-context capabilities of jina-embeddings-v3 and multilingual-e5-large, limiting their applicability in complex code search scenarios.
III-C Test-Driven Agents
Traditional test generation approaches rely heavily on human expertise, but recent work demonstrates the potential of automated test agents. LLM-based agents have shown promising results in various testing scenarios. For instance, in the domain of deep learning libraries, agents capable of generating diverse and meaningful test inputs are proposed [46, 47], alongside other studies exploring LLM-based agents for specialized tasks. For test oracle generation, few-shot prompting is used to generate context-aware assertions for test oracle generation [48, 49], achieving a high accuracy. In program repair, some studies [50, 51, 52, 53, 54] demonstrate the effectiveness of integrating LLMs with iterative validation workflows [55]. However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance [56].
Figure 2: The construction process of CoSQA+. (a) Queries sourced from CoSQA and code snippets filtered from CodeSearchNet. (b) Candidate pairs selected based on query-code cosine similarity, computed using a multi-model approach. (c) Pipeline for annotation by test-driven agents.
IV CoSQA+
Fig. 2 illustrates the benchmark construction process for CoSQA+. The process starts with query and code collection, in which we leverage the queries from CoSQA and collect python code snippets filtered from CodeSearchNet. In candidate pairs construction, collected code snippets and queries are embedded as vectors to calculate similarity. For each query, the top 20 query-code pairs with the highest similarity are selected as candidates for potential matches. In model annotation, each query-code pair is annotated by test-driven agents with “1” or “0”, indicating whether the code snippet exactly matches the query. These labeled pairs are then added to CoSQA+.
IV-A Query and Code Collection
This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus [7] serves as the code sources.
Query Source
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes [8]. We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets [57]. (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent. This careful curation helps maintain high data quality and reduces noise in the training process [58]. By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
Code Source
To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages [7]. According to our survey, code searchers prefer function codes with clear input and output. These codes are high quality with higher readability, testability and maintainability [59]. To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach. This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement. Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
IV-B Candidate Pairs Construction
Labeling all query-code pairs is impractical. To improve efficiency, we only retain high-confidence pairs before annotation. This section details the construction process of high-confidence query-code pairs.
Construction Steps
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. Fig.  3 shows a detailed breakdown of query distributions by the number of exactly matched codes.
Figure 3: Query counts grouped by number of matched codes
Multi-model Embedding
High similarity between query and code is fundamental for a high-quality dataset [57]. During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. This potentially leads to unfair test results of the benchmark. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. We empirically validate this practice. Specifically, we use a single model to select 20 candidate codes for each query and use an ensemble of models, CodeBERT [12], UniXcoder [31] and CodeT5+ 110M embedding [34], to select another 20 candidates. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. We calculate the overlap between the candidates chosen by the single model and those chosen by the ensemble and analyze the overlap between them. Although bias cannot be completely eliminated, the experiment result in Table I shows that comprehensively utilizing multiple models for matching can effectively reduce the bias of query-code pairs toward any single model. As can be seen, the proportion of overlapping codes is low, indicating that the codes matched by multiple models for a query do not show a significant preference for any particular single model.
TABLE I: Overlap ratio of candidate codes between individual models and the ensemble of models.
Model Overlap Ratio
CodeBERT 0.135
UniXcoder 0.164
CodeT5+ embedding 0.088
The relatedness between the query and code is quantified by the cosine similarity, calculated as:
${cosine\_similarity}=\frac{\mathbf{v}_{\text{query}}\cdot\mathbf{v}_{\text{code}}}{\|\mathbf{v}_{\text{query}}\|\|\mathbf{v}_{\text{code}}\|}$
where $\mathbf{v}_{\text{query}}$ and $\mathbf{v}_{\text{code}}$ are the embedding vectors of the query and code, respectively.
By taking the average similarity score from these three models, we achieve a more balanced and fair assessment of query-code pair quality.
Top-20 Code Candidates Selection
Before constructing the full set of candidate pairs, we randomly select 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query. The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Fig. 4). As shown in our analysis, when the number of candidates reaches 20, the marginal benefit of further expansion diminishes. This decline can be attributed to limitations of codebase, nature of queries and ability of embedding models. Furthermore, adding just one more candidate code per query results in an additional 20,604 pairs, which would incur substantial costs in terms of both time and financial resources for annotation. Thus, limiting candidates to the top-20 strikes a balance between maximizing the retrieval of relevant codes and controlling annotation costs.
Figure 4: Probability of identifying matched codes as the number of candidate codes increases
IV-C Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design the test-driven annotation agents based on DeepSeek-V3 [60], using the algorithm shown in Algorithm 1. Initially, A preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation.
Algorithm 1 Test-Driven Agent Annotation
Input: query $q$, code snippet $c$, max_try_time
Output: label $l\in\{0,1\}$
1:  $assessment\leftarrow\text{preliminary\_screener}(q,c)$
[…]
3:     return  1
[…]
5:     return  0
[…]
7:     $test\_prog\leftarrow\text{generate\_test\_program}(q,c)$
8:     $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
9:     $try\_time\leftarrow 0$
10:     while $exec\_res.\text{error}\ \textbf{and}{}\ try\_time\leq max\_try\_time$ do
11:        $terminal\_commands\leftarrow\text{fix\_test\_program}(test\_prog,exec\_res.\text{error})$
[…]
13:        $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
14:        $try\_time\leftarrow try\_time+1$
15:     end while
[…]
17:  end if
Figure 5: Example illustration of (1) clear match, (2) clear non-match, and (3) unclear cases.
Preliminary Screener
The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Fig. 5. Guided by the prompt instructions, the model assesses whether the code fully addresses the query’s functional requirements. For unambiguous cases, it provides a justification and immediate classification. For ambiguous cases (representing 33% of candidates), it flags the pair for deeper verification by downstream agents. This triage mechanism optimizes the annotation workflow by resolving straightforward cases upfront while reserving computational resources for ambiguous scenarios.
Test Program Generator
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, we generate the test program based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation.
Test Executor and Bug Fixer
The test executor operates within an isolated Docker environment to ensure safe execution while preventing host system interference. When errors such as ModuleNotFoundError occur, the pipeline initiates a two-stage repair process. First, it attempts direct installation of missing dependencies using standard package managers. If unresolved, the system invokes an LLM to generate context-aware terminal commands for dependency resolution, which are executed iteratively until the error is mitigated or deemed irreparable. After dependency fixes, the test suite is re-executed to validate functionality. Execution traces, assertion outcomes, and error logs are aggregated into a structured report, which is forwarded to the final arbiter for assessment. This approach enhances the reliable verification of ambiguous cases while maintaining environmental integrity, achieving an executable program ratio of 83.67%.
Final Arbiter
The framework ultimately concludes with a final arbiter stage, where the agent employs a specialized prompt to assess if the generated code fully satisfies the query. This assessment is based on both the test program produced and the outcomes of its execution. This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior rather than relying exclusively on static structural analysis and semantic analysis. Fig. 6 illustrates the process of handling an unclear case.
Figure 6: An example for handling an unclear case through test program generation, test execution, and final arbitration.
IV-D Benchmark Statistics
In total, 412,080 labeled query-code pairs are retained. Table II shows the statistics of our CoSQA+ benchmark.
TABLE II: Statistics of CoSQA+. "#" denotes the count, "Code avg. len." is the average number of characters per code snippet, and "Exec. rate" represents the percentage of test programs that executed successfully.
# of pairs # of code Code avg. len. Avg. test cases Exec. rate
412,080 132,952 378.052 4.31 83.67%
IV-E Cost Analysis
For data annotation, a total of 412,080 data samples are labeled by Deepseek-V3 at a cost of approximately \$257, resulting in an average cost of \$0.00062 per sample. The labeling time varies significantly based on complexity: clear cases require approximately 10 seconds per sample, while non-clear cases may take up to 50 seconds, depending on API rate limits, dependency installations, and network latency. However, the annotation process benefits from full concurrency across multiple queries, enabling rapid completion when computational resources are sufficient.
In contrast, manual annotation by Python experts involves substantial time investment. On average, a Python expert spends 8 minutes per sample due to extensive research and analysis. For test-driven annotation, where experts design tailored test programs for code-query pairs, the time increases to 34 minutes per sample, with complex functions exceeding one hour.
IV-F Informed consent
This study involved human participants. Informed consent was obtained from all participants prior to their involvement in the study. Participants were informed about the purpose of the research, the procedures involved, the potential risks and benefits, and their right to withdraw from the study at any time without penalty. All data collected were anonymized to ensure participant confidentiality.
V Evaluation
In this section, we conduct experiments with corresponding analysis to answer the following research questions (RQs):
RQ1: Is the quality of CoSQA+ higher than that of CoSQA?
RQ2: Can our test-driven agents surpass Python experts and LLMs in the annotation task?
RQ3: How effective are the existing approaches in multi-choice code search on our benchmark?
RQ4: How does each agent contribute to the performance of test-driven agents?
V-A RQ1: Quality Comparison between CoSQA+ and CoSQA
We design two experiments to verify whether CoSQA+ offers superior code quality and training performance compared to CoSQA.
Metrics
Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG [14] and MRR [12]. NDCG measures the quality of ranking by considering both the relevance and position of retrieved results. For a given query $q_{i}$, we calculate NDCG@k as:
$NDCG@k=\frac{DCG@k}{IDCG@k}$
where DCG@k is the Discounted Cumulative Gain up to position k, and IDCG@k is the ideal DCG@k (the maximum possible DCG@k). The final NDCG score is averaged across all queries. MRR evaluates the average reciprocal of the rank at which the first relevant code appears:
$MRR=\frac{1}{|N|}\sum_{i=1}^{|N|}\frac{1}{rank_{i}}$
where |N| is the number of queries and rank_i is the rank of the first relevant code for the i-th query.
Experimental Setup
In the first experiment, we randomly select 1000 query-code pairs labeled "1" from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, from which we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
For the second experiment, We fine-tune CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ and CoSQA separately, testing the models on the CSN Python with NDCG@10 and MRR as evaluation metrics. From the CSN benchmark we use for validation, known as CSN99, we extract Python test samples and build 399 query-code pairs for 95 queries, referred to as CSN Python. The queries in CSN99 are realistic queries sourced from Bing, carefully selected and verified by humans. In our training process, the model is trained using individual query-code pairs, similar to traditional one-to-one code search models. The CoSQA+ dataset is divided into training, validation, and test sets in an 8:1:1 ratio, while CoSQA is split according to its original division.
TABLE III: Performance of code search models fine-tuned with different data. The test set is CSN Python.
Model Metric Fine-tune Dataset
None CoSQA CoSQA+
CodeBERT MRR 0.952 0.939 0.966
NDCG@10 0.958 0.937 0.962
UniXcoder MRR 0.891 0.951 0.954
NDCG@10 0.890 0.946 0.952
CodeT5+ MRR 0.794 0.938 0.940
embedding NDCG@10 0.806 0.943 0.951
Results and Analysis
After the volunteers compare the two datasets, we come to the result that out of the 2000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table III, the three models fine-tuned with CoSQA+ outperform those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
RQ1 Summary: CoSQA+ has better quality than CoSQA. Fine-tuning CodeBERT, UniXcoder and CodeT5+ embedding with CoSQA+ yields better performance than with CoSQA, especially for multiple relevant codes.
V-B RQ2: Performance of Agent-based Annotator
We design this experiment to assess whether our agent-based annotation approach can achieve comparable or superior quality to traditional human annotation and single LLM annotation.
Experimental Setup
To establish a reliable ground truth, we first have three Python experts design test programs for 400 randomly selected query-code pairs and annotate them based on test execution results. The final ground truth labels are determined by majority voting among these three experts. We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [61].
Results and Analysis
TABLE IV: Accuracy comparison of different annotation methods against ground truth.
Method Accuracy
Agent-based Annotator 0.920
Human Wo. Test Program 0.880
Deepseek-V3 0.865
O1-mini 0.833
Claude-3.5-Sonnet 0.845
As shown in Table IV, our agent-based annotator achieves 92.0% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (84.0%), O1-mini (82.0%) and Claude-3.5-Sonnet (84.0%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
Importantly, about 83% of the test programs generated by our agent are successfully executable, indicating the reliability of our test program generation approach. The high executability rate indicates that most query-code pairs can be verified through actual program testing rather than relying solely on semantic analysis.
The superior performance of our agent-based annotators can be attributed to two key factors: 1) The test program provides concrete evidence of whether the code satisfies the query requirements, reducing ambiguity in the annotation process; 2) The systematic verification through test execution helps eliminate human bias and inconsistency that often occur in manual annotation.
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.818, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 92.0% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
V-C RQ3: Existing Methods Performance on CoSQA+
This experiment aims to evaluate the effectiveness of different models in accurately retrieving code snippets that match natural language queries on the CoSQA+ benchmark.
Task
The natural language code search task is a text retrieval problem. Given a natural language query $q_{i}$ and a collection of code snippets $C=\{c_{1},\ldots,c_{n}\}$, the goal is to identify all code snippets
$\{c^{*}_{j}\mid c^{*}_{j}\in C\text{ and }c^{*}_{j}\text{ exactly matches }q_{i}\}$
that exactly match the query.
Experimental Setup
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric. NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly. To further comprehensively evaluate the performance of the models, we also introduce Mean Average Precision (MAP) and Recall as additional metrics. MAP provides a measure of the overall ranking quality by averaging the precision scores across all relevant results, which is especially useful for scenarios where the order of multiple correct matches matters. Meanwhile, Recall evaluates the model’s ability to retrieve all relevant code snippets from the dataset, ensuring that no important results are missed. Together, these metrics offer a more robust assessment of the model’s effectiveness in semantic code search tasks. The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach [27] with state-of-the-art semantic code search models, namely: CodeBERT [12], UniXcoder [31], and CodeT5+ embedding [34]. In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3 [35], All-MiniLM-L12-v2 [40], All-mpnet-base-v2 [41], and Multilingual-e5-large [38].
Results and Analysis
TABLE V: Performance of Various Methods on CoSQA+ Using NDCG@10, MRR, MAP and Recall Metrics.
Method NDCG@10 MRR MAP Recall
BoW 0.037 0.027 0.027 0.016
CodeBERT 0.232 0.175 0.164 0.149
UniXcoder 0.187 0.140 0.132 0.110
CodeT5+ embedding 0.179 0.136 0.127 0.101
Jina-embedding-v3 0.225 0.171 0.160 0.145
All-MiniLM-L12-v2 0.214 0.163 0.153 0.135
All-mpnet-base-v2 0.227 0.172 0.162 0.145
Multilingual-e5-large 0.218 0.166 0.156 0.136
While traditional information retrieval methods, such as BoW, are computationally efficient, Table V demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist. Among the deep learning-based models, CodeBERT achieves the highest performance with NDCG@10 of 0.232, MRR of 0.175, MAP of 0.164 and Recall of 0.149. Other embedding-based models also show substantial improvements over traditional methods, highlighting the benefits of modern deep learning architectures for semantic code search. This strong performance can be attributed to their encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances their code understanding capabilities. The embedding models originate from different time periods and vary in parameter sizes. Surprisingly, CodeBERT outperforms the latest embedding models, which may be attributed to the fact that CodeBERT tested here is fine-tuned specifically for code search tasks, whereas the latest embedding models are designed for general-purpose tasks. Some smaller-parameter embedding models demonstrate impressive performance. For instance, all-mpnet-base-v2, with only 109M parameters, and all-MiniLM-L12-v2, with a mere 33M parameters, still exhibit remarkable performance, surpassing many models with larger parameter sizes. This represents a significant advancement in embedding models in recent years, providing a solid foundation for the practical implementation of semantic search based on large models.
From the evaluation results, the performance differences among embedding models seem not significant. However, due to the richness of CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models. Moreover, embedding models generally perform poorly on our benchmark. From the perspective of model design, embedding models are mainly used for retrieval tasks. To improve inference speed, their weight parameters are usually small, limiting their ability to deeply understand query sentences and code.
The CoSQA+ is challenging in two main aspects: First, the queries originate from CoSQA, which are real user queries filtered from the Bing search engine. These queries are often expressed casually and may contain spelling or grammar errors. Second, the codebase used for retrieval is not only large in scale but also contains many confusing code snippets.
RQ3 Summary: The evaluation on CoSQA+ clearly indicates that deep learning-based models significantly outperform traditional retrieval methods for semantic code search. Notably, CodeBERT leads the performance ranking by achieving the highest scores in both NDCG@10 and MRR, while the alternative embedding-based models also offer competitive results.
V-D RQ4: Contributions of Different Components in Test-Driven Agents
This experiment is to assess the contribution of each component in our test-driven agent framework.
Experimental setup
The final test-driven agents comprise the following components: Preliminary Screener (PS), Test Program Generator (TPG), Test Executor (TE), Bug Fixer (BF), and Final Arbiter (FA). We divided the agent into different configurations by combining these components in various ways. For each configuration, we compared the predicted labels with the ground truth labels (obtained from human evaluation based on the test programs) to calculate the accuracy. Additionally, we compare the Simulated Executor with both the Test Executor and the Bug Fixer to evaluate whether the large language model can accurately simulate execution without actual interpretation. Unlike the Test Executor, which runs the test program using a Python interpreter, the Simulated Executor leverages the LLM’s reasoning capabilities to emulate execution and directly predict the outcome.
Result and Analysis
Table VI shows the accuracy for different configurations.
Configuration Accuracy
Single LLM 0.865
PS + FA 0.840
PS + TPG + FA 0.833
PS + TPG + TE + FA 0.835
PS + TPG + Simulated Executor + FA 0.840
PS + TPG + TE + BF + FA 0.920
TABLE VI: Accuracy of Various Test-Driven Agent Configurations (Abbreviated)
The results indicate that the baseline Single LLM achieves an accuracy of 0.865. When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.840. Incorporating the Test Program Generator alongside these yields an accuracy of 0.833, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [62]. Adding the Test Executor to the configuration improves the performance marginally (0.835), and using a Simulated Test Executor does not provide a significant advantage (0.840). This aligns with the observation that simulated execution cannot fully capture the complexity of the real compile-and-run process, potentially introducing inaccuracies or misleading feedback. Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.920. This enhancement demonstrates that the Bug Fixer is crucial for processing the test programs effectively, as it enables the system to generate real execution results that are vital for improving prediction accuracy.
RQ4 Summary: The Bug Fixer is the most important component in the test-driven agent framework as it makes the test program executable and yields effective test results. However, it relies on the preceding steps, such as generating the test program. Simulated execution results cannot replace the benefits of real execution feedback.
VI Discussion
VI-A Test Misjudgment
Previous studies show that as the size of the codebase increases, the challenges faced by embedding models also increase [63, 64]. This conclusion aligns with intuition, but with the analysis of other datasets and the construction of CoSQA+, we find that the decline in performance is not only due to the need to exclude more confusing codes but also because of the increased likelihood of test misjudgments. For example, a LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set. As a result, this correct selection is misjudged as incorrect.
When constructing code search datasets, query sets and code sets are typically extracted from original query-code pairs for test. However, the number of all possible query-code pairs (similar to a Cartesian product) formed between the query set and the code set far exceeds the number of original pairs. Existing methods assume that the code of a pair outside the original pairs cannot correctly match the query, but this assumption is not entirely valid. As shown in Fig. 7, in CoSQA+, excluding codes without satisfied query, 36.7% of codes can satisfy two or more queries. Constructing a dataset completely free of misjudgments is almost impossible, as it would require checking all possible pairs. As can be seen by combining Fig 3 and Fig. 7, the approach adopted in CoSQA+, where 20 candidate codes are annotated for each query, helps mitigate these testing misjudgments. Another possible measure is to invoke test-driven annotation agents when a code search method selects a previously unseen query-code pair to determine whether the selected code matches the query. However, this approach may introduce significant computational overhead, especially when the candidate code repository is very large.
Figure 7: Distribution of codes by the number of queries they satisfy
VI-B Inherent Ambiguity in Query and Code Labels
A major challenge in code search is the inherent ambiguity in queries and code. Queries are often vague or poorly phrased, leading to multiple interpretations, while code snippets may be incomplete or abstracted, making correctness subjective. For instance, a query like "how to check if a folder exists and create it if not" can be interpreted differently based on assumptions about paths or edge cases. Similarly, code may solve the problem partially or in a non-standard way, complicating strict labeling ("exact match"). Test-driven annotation helps reduce ambiguity by generating executable test programs aligned with the query’s intent. However, variability in query interpretation and incomplete code still introduce labeling challenges. Future improvements will require clearer queries and more complete code snippets to minimize ambiguity.
VI-C Ethical Consideration
In this work, we prioritize ethical considerations by the responsible use of data and methodologies. The dataset sources, CoSQA and CodeSearchNet, are publicly available under permissible licenses, mitigating privacy concerns. Our test-driven agents are designed to minimize human bias by automating code evaluation, thus enhancing fairness and consistency. Additionally, we openly release CoSQA+ to promote transparency and reproducibility in code search research. While our approach aims to improve code search efficiency, we acknowledge the potential for misuse in generating or retrieving malicious code. We encourage the community to adopt ethical guidelines when using our benchmark and tools, ensuring that their application aligns with positive social impact.
VII Conclusion
In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA. The test-driven annotation approach achieves a high accuracy of 92.0%, outperforming traditional human annotation and single LLM methods. Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks. We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task.